OpenStudy (anonymous):

What is the solution to the system of equations? 5x+4y=8 x-2y=10 please help :) will fan and medal

OpenStudy (alexandervonhumboldt2):

x=10+2y

OpenStudy (alexandervonhumboldt2):

5*(10+2y)+4y=8

OpenStudy (alexandervonhumboldt2):

50+14y=8

OpenStudy (alexandervonhumboldt2):

y=-3 x=4

OpenStudy (anonymous):

thank you :)